OpenStudy (anonymous):

What is the value of 2x+1/2 when 5x-3=-3x+5?

OpenStudy (anonymous):

5x-3=-3x+5 5x + 3x = 5 + 3 8x = 8 x = 1 So 2x+1/2 = 2(1) + 1/2 = 2+1/2 = 5/2

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

yes thank youuu